How do i solve -1 2/7m=6?
Answer by sdmmadam@yahoo.com(530)

How do i solve -1 2/7m=6?
-1 2/7m=6?
(-12) = 6X(7m) (multiplying through out my 6m)
Dividing by (6X7)
(-12)/(6X7) = m
(-2/7) = m (cancelling 6 in the nr and in the dr)
Answer: m = (-2/7)